# Pretraining-finetuning Framework for Efficient Co-design: A Case Study on Quadruped Robot Parkour

### v1

@@ II-C1 Fine-tuning Based on Pre-trained Model
By integrating spatial domain randomization with regularization, we have developed a control strategy applicable to robots with varying leg lengths. However, this strategy does not guarantee optimality in each morphology. To provide a more reliable fitness function for the morphology optimization algorithm, each morphological candidate provided by the morphology optimization algorithm is fine-tuned. Diverging from the training of generalized models, we made three modifications: (1) The spatial domain randomization is omitted, focusing the training exclusively on the specific morphology. (2) Remove the regularizations. (3) The difficulty level of the curriculum training is elevated to the highest tier. Through these modifications, the fine-tuned control strategy is better suited for robots in the specific morphology.
@@ II-C2 Morphology Optimization
The objective of morphology optimization is to find optimal morphology parameters $\xi\in\Xi$ for a specified parkour task $\kappa\in\mathcal{K}$ and a given control strategy $\pi\in\Pi$, such that the objective function $f(\left.\xi\right|\kappa,\pi):\Xi\times\mathcal{K}\times\Pi\to\mathbb{R}$ is maximized. This study focuses primarily on the high jump and long jump tasks, denoted as $\mathcal{K}=\left\{{{\kappa_{high}},{\kappa_{long}}}\right\}$. The optimization objective (also known as fitness) $f(\cdot)$ can be any metric, we use the robot’s non-discounted cumulative reward in a single epoch as our optimization objective, as shown below:
$f=\frac{1}{N}\sum\limits_{i=0}^{N}{\sum\limits_{t=0}^{T}{r(s_{t}^{i},a_{t}^{i})}}$ (3)
where $T$ denotes the number of steps in an episode, while $N$ represents the number of robots interacting with the environment. Typically, $\xi$ is non-differentiable relative to $f(\cdot)$ because legged locomotion involves many discrete changes due to foot contact. Hence, we utilize a black-box optimization method, specifically BO, for the optimization of morphology parameters. The optimization problem can be summarized as follows:
@@ III-A Implementation Details
During the training process, we employed Isaac Gym as the simulator. The network was built using PyTorch, and training was facilitated using Nvidia 3090 GPUs. The neural network policy generated PD targets for all motors, operating at a frequency of 50 $\mathrm{H}\mathrm{z}$. Concurrently, the PD controllers functioned at a higher frequency of 200 $\mathrm{H}\mathrm{z}$, with a $k_{p}$ set to 40 and a $k_{d}$ configured to 0.7.
During the training process, we set the number of robots (represented as $N$) to 6144. This also means that our spatial domain randomization method created $N$ different robots with varying leg lengths. Specifically, as shown in Section II-B1, during the initialization of the environment, we first select $\xi_{i,i=0,1,2,3}$ within the range $\mathcal{U}(0.6,1.4)$. After obtaining $N$ sets of $[\xi_{0},\xi_{1},\xi_{2},\xi_{3}]$ parameters, we update the original URDF files based on Tab. I to obtain $N$ sets of different URDF files, which are then loaded into the simulator.
III-B Performane of Control Policy
Comparison of Domain Randomization: To evaluate the effectiveness of spatial domain randomization (DR), we conducted three experiments for comparison.
  • Temporal-DR. We varied the leg length of the robots during the training process. Since altering the leg lengths of the robot in each episode is impractical, we selected eight typical morphologies. We start training with the first set of morphologies, switching to the next set once $1/8$ of the overall process is completed, and so on, until the training concludes.
  • Spatial-DR (Ours). The proposed method, involved using $N$ robots with varied morphologies to simultaneously gather data for training the policy.
  • No-DR. No domain randomization strategy was employed, and all robots shared the same default morphology, i.e., $\xi_{i}=[1.0]\times 4$.
To validate the performance of three domain randomization methods, we uniformly sampled dozens of morphologies and recorded the average cumulative rewards achieved by each method under these morphologies. Furthermore, we conducted T-tests comparing the proposed method with Temporal-DR and No-DR, and the results are presented in Tab. II. It is evident that the proposed method yielded the highest mean values, and the P-values between the proposed method and the two baselines are less than the threshold of 0.05, confirming the statistical superiority over the two baselines. These findings demonstrate that the Spatial-DR method effectively enhances the generality of the control policy without necessitating extra computational resources.
@@ III-B Performane of Control Policy
Mean value 5.11$\pm$0.02 10.75$\pm$0.02 7.50$\pm$0.01
P-value $1.18\times 10^{-9}~{}~{}~{}~{}~{}8.36\times 10^{-9}$
Comparison of Regularization: To verify the effectiveness of regularization methods, we designed the following experiments:
  • Activation Regularization (A-Reg). As demonstrated by Eq. (1), we added a regularization term to the Critic’s loss function. Following numerous tests, we selected a coefficient $\beta=0.2$.
  • Discount Regularization (D-Reg). As detailed by Eq. (2), we altered the discount factor $\gamma$ in the PPO algorithm to a regulated value $\gamma_{\rm reg}$. After numerous tests, we settled on $\gamma_{\rm reg}=0.98$.
  • Normal. This approach does not employ regularization, retaining the original $\gamma$ value of 0.99.
  • Phasic Policy Gradient (PPG) [27]. A SOTA method for enhancing the generalizability of RL algorithms. It integrates auxiliary losses during the training of the PPO algorithm to improve the algorithm’s generalizability, details can be found in [27].
@@
Figure 4: The curve of cumulative rewards changes with the training steps. The results presented are average across three experiments, and the shaded areas represent the variance of these experiments.
@@
To compare the effectiveness of different regularization methods, we plotted the cumulative reward curves for each approach during the training process, as shown in Fig. 4. It is evident that D-Reg yielded the best results, followed by A-Reg, while PPG underperformed, even falling behind Normal, which doesn’t use any regularization techniques. Taking both the effectiveness and the simplicity of algorithm implementation into account, D-Reg proves to be superior in our locomotion control tasks. It is noteworthy that PPG exhibited the poorest performance. We suspect this may be because the PPG algorithm was developed for image-based RL and might not be well suited for motion control RL with low dimensional observations.
Performance of Fine-tuning: In this section, we demonstrate the effectiveness of fine-tuning, considering both the time consumption and the performance enhancements, we applied two methods as follows.
  • Standard Training. Normative training under the specific morphology without pre-training.
  • Fine-tuning. The proposed method, starts with pre-training using the spatial domain randomization and regularization technique, followed by fine-tuning under the specific morphology.
Figure 5: Heatmap of co-design experiments. Each block represents the cumulative rewards achieved using the corresponding control strategy in the parkour task under a given morphology. The $x$-axis represents the ratio factor between the thigh and the shank of the front leg, while the $y$-axis indicates the ratio factor between the thigh and shank of the hind leg. The darker the shade of each block, the higher the cumulative rewards, indicating better performance of the algorithm. For Online-PCODP, Offline-PCODP, and EAT, the eight typical morphologies used are the same as Fig. 6(b).
Figure 6: Fine-tuning experimental results. (a) Time consumption. The subgraph on the right is an enlarged version of the Fine-tuning curves. Each process was run three times, the solid lines represent the mean value, while the shaded areas indicate the variance. (b) shows the bar chart of cumulative rewards on eight typical morphologies before and after fine-tuning, displaying the mean results of three runs, the $y$-axis represents the eight typical morphologies.
@@
The time comparison, illustrated in Fig. 6(a), reveals that Standard Training begins to converge at the 2000th epoch, while Fine-tuning starts converging at just the 100th epoch. These findings validate that the proposed method can significantly accelerate the training process. For instance, when training 30 specific morphologies, Standard Training requires $30\times 6000=1.8\times 10^{5}$ epochs, whereas Fine-tuning only needs $6000+30\times 400=1.8\times 10^{4}$ epochs. The number 6000 represents the pre-training steps, while 400 indicates the fine-tuning steps tailored for specific morphologies, equivalent to 6.67% of the former. Furthermore, we also compared the performance enhancements. Fig. 6(b) displays the accumulative rewards in eight typical morphologies, both using the pre-trained model (Before FT) and after fine-tuning the pre-trained model (After FT). It indicates that, in all cases, the performance of the strategies improved significantly after fine-tuning, thereby affirming the effectiveness of our proposed method in terms of performance enhancement.
@@ III-C Comparison with Co-design Baselines
In this section, we compare the proposed method with three co-design baselines, we first introduce these methods.
  • Online Policy Conditioned on Design Parameters (Online-PCODP) [6]. This method treats design parameters as privileged information and samples these parameters uniformly within their range during the training process. It assumes that near-optimal performance can be achieved using a policy conditioned on design parameters. In our parkour task, since modifying morphological parameters in each episode is impractical, we have adopted the Temporal-DR method to train the policy.
  • Offline Policy Conditioned on Design Parameters (Offline-PCODP) [3]. This method collects interaction data from various morphologies using online RL and then uses offline RL to create a general policy from these data, which acts as a surrogate for optimizing the morphology. In detail, it first trains separate policies for eight typical morphological parameters with PPO, and then develops the general policy using TD3-BC [30].
  • Embodiment-aware Transformer (EAT) [5]. Similar to Offline-PCODP, this approach initially focuses on training separate control policies for various morphological parameters to gather interaction data. These data are then trained using a network based on the Transformer architecture, ultimately yielding a control policy with generalization capabilities.
  • Pre-training and Fine-tuning (Ours). The proposed method. We start by training an initial model with generalization, followed by fine-tuning for different morphologies.
@@
We uniformly sampled 81 different morphologies from the sampling range and plotted the cumulative rewards of different methods under these morphologies in Fig. 5. As can be seen from Fig. 5(d), our proposed method achieves the highest rewards across various morphologies, demonstrating that the pre-training and fine-tuning approach can provide a more effective control strategy for the optimization of morphological parameters. For the Online-PCODP method, as observed in Fig. 5(a), its highest reward values are concentrated under the last trained morphology. For previously encountered morphologies such as the extremely short legs and extremely long legs, it shows lower cumulative rewards. These results indicate that this method is prone to forgetting previously trained morphologies, focusing instead on the final morphology. For the Offline-PCODP method, as seen in Fig. 5(b), its highest rewards are found near the eight typical morphologies encountered during training. However, for morphologies not previously encountered, it shows lower cumulative rewards, proving that this method struggles to generalize to morphologies significantly different from those seen during training. Regarding the EAT method, as seen in Fig. 5(c), its performance appears to be poor across all morphologies. We speculate that this may be due to the original EAT method focusing on simple planar walking tasks with low-dimensional states such as body linear velocity and angular velocity. Our method, on the other hand, addresses complex parkour tasks, where terrain information is crucial but results in higher state dimensionality, potentially causing Transformer-based algorithms to fail.
Furthermore, we applied the aforementioned methods for co-design in both the high jump and long jump tasks. We recorded the accumulative rewards of the optimal morphologies and their corresponding control strategies obtained by each method within a single episode, as demonstrated in Tab. III. It is evident that the proposed method achieved the highest values in both tasks, markedly surpassing the baseline methods. This substantiates the effectiveness of the proposed pretraining-finetuning framework in the context of co-design. In Tab. III, we also recorded the time consumption of various methods, wherein the time for the Online-PCODP and the proposed methods encompasses the duration for training the control strategy as well as the time for BO. The time for Offline-PCODP and EAT contains the duration for training the expert control strategy, collecting expert data, training the general network, and the time for BO. The introduction of the fine-tuning during the BO process has incurred some additional time, however, overall, the time consumed by the proposed method is still much less than that of the Offline-PCODP and EAT methods. Taking both performance and time consumption into consideration, the advantages of the proposed method are clearly notable.
@@ III-D Parkour Performance Comparison
Comparison of Single Task: In this section, we compared the performance achievable by baselines with that of the proposed approach for the tasks of long jump and high jump, as illustrated in Tab. IV. The values for the baseline methods were cited from the corresponding papers. As seen in Tab. IV, the proposed method significantly enhances the performance of robots in various parkour tasks, demonstrating that, in addition to improving the motion controller’s performance, appropriate optimization of the mechanical structure can also enhance the performance of quadruped robots in extreme tasks. Furthermore, we illustrated the performance of the robots in extreme tasks before and after morphological optimization, as shown in Fig. 7. For the long jump task, it is clear that the optimized robots are equipped with longer hind legs, which can provide more robust propulsion, while shorter front legs can assist the robot in gripping the ground ahead to prevent falls. Similarly, for the high jump task, the robots evolved to have longer legs, aiding in climbing. Both experiments underscore the potential of the co-design approach in unlocking the parkour performance of robots.
Comparison of Mixed Task: In the previous section, we conducted co-design of the robot for different tasks. In this part, we perform co-design of the robot for the mixed tasks and compare it to quadruped mammals. The crural index, the ratio of tibia (the hind shank) to femur (the hind thigh) length times 100, is higher in good jumpers, with an average crural index of 80 in humans and typically exceeding 100 in animals proficient in jumping [31]. We measured the fitness of morphologies with crural indices above and below 100 during BO, using Levene’s test and T-test for analysis. The data in Tab. V show that morphologies with higher crural indices performed better, in line with natural findings. The Levene test checks if data sets have similar variances, a value above 0.05 confirms this. We use the T-test, which assumes homogeneity of variances, to analyze the two sets of data. A P-value below 0.05 shows they are statistically different.
@@
TABLE V: Fitness of different crural index
Crural Index Levene-test P-value T-test P-value
>100 $\leq$ 100
29.177 $\pm$ 2.02 27.898 $\pm$ 1.76 0.608 0.027

### v2

@@ II-C1 Fine-tuning Based on Pre-trained Model
By integrating spatial domain randomization with regularization, we have developed a control strategy applicable to robots with varying leg lengths. However, this strategy does not guarantee optimality in each morphology. To provide a more reliable fitness function for the morphology optimization algorithm, each morphological candidate is fine-tuned. Unlike the training of generalization models, we removed the spatial domain randomization method during the fine-tuning process. Instead, we created thousands of identical robots to interact with the environment, allowing us to fine-tune specifically for the targeted morphology.
@@ II-C2 Morphology Optimization
The objective of morphology optimization is to find optimal morphology parameters $\xi\in\Xi$ for a specified parkour task $\kappa\in\mathcal{K}$ and a given control strategy $\pi\in\Pi$, such that the objective function $f(\left.\xi\right|\kappa,\pi):\Xi\times\mathcal{K}\times\Pi\to\mathbb{R}$ is maximized. This study focuses primarily on the high jump and long jump tasks, denoted as $\mathcal{K}=\left\{{{\kappa_{high}},{\kappa_{long}}}\right\}$. The optimization objective (or fitness) $f(\cdot)$ can be any metric, we use the robot’s non-discounted cumulative reward in a single epoch as our optimization objective, as shown below:
$f=\frac{1}{N}\sum\nolimits_{i=0}^{N}{\sum\nolimits_{t=0}^{T}{r\left({s_{t}^{i},a_{t}^{i}}\right)}}$ (4)
where $T$ denotes the number of steps in an episode, while $N$ represents the number of robots interacting with the environment. In our previous work [3], we demonstrated that the BO method is well-suited for optimizing morphological parameters in co-design, given that $\xi$ is non-differentiable with respect to $f(\cdot)$. We can summarize the optimization problem as follows:
@@ III-A Implementation Details
During the training process, we employed Isaac Gym as the simulator. The network was built using PyTorch, and training was facilitated using Nvidia 3090 GPUs. The neural network policy generated PD targets for all motors, operating at a frequency of 50 $\mathrm{H}\mathrm{z}$. Concurrently, the PD controllers functioned at a higher frequency of 200 $\mathrm{H}\mathrm{z}$. During the training process, we set the number of robots (represented as $N$) to 6144. Specifically, as shown in Section II-B1, during the initialization of the environment, we first select $\xi_{i,i=0,1,2,3}$ within the range $\mathcal{U}(0.6,1.4)$. After obtaining $N$ sets of $[\xi_{0},\xi_{1},\xi_{2},\xi_{3}]$ parameters, we update the original URDF files based on Tab. I to obtain $N$ sets of different robots, which are then loaded into the simulator.
@@
Mean value 3.51$\pm$0.30 15.58$\pm$0.67 9.91$\pm$0.37
P-value $4.42\times 10^{-5}~{}~{}~{}~{}~{}4.55\times 10^{-4}$
III-B Performane of Control Policy
Comparison of Domain Randomization: We employed the following three methods to demonstrate the effectiveness of spatial domain randomization (DR). Temporal-DR. This method varied the leg length of the robots during the training process. Specifically, it uniformly sampled 81 morphologies from the sampling range (as shown in Fig. 6). It started the training with the first set of morphologies, then switched to the next set once 1/81 of the overall process was completed, and continued this pattern until the training concluded. Spatial-DR (Ours). The proposed method, involved using $N$ robots with varied morphologies to simultaneously gather data for training the policy. No-DR. No domain randomization strategy was employed, and all robots shared the same default morphology, i.e., $\xi_{i}=[1.0]\times 4$. We randomly sampled one hundred morphologies and recorded the average cumulative rewards under these morphologies. Furthermore, we conducted T-tests comparing the proposed method with Temporal-DR and No-DR, and the results are presented in Tab. II. It is evident that the proposed method yielded the highest mean values, and the P-values between the proposed method and the two baselines are less than the threshold of 0.05, confirming the statistical superiority over the two baselines. In addition, we measured the time required by each method. The time taken for Spatial-DR is approximately the same as that for No-DR. In contrast, Temporal-DR is slower, as it requires continuous termination of the current environment and the loading of new morphologies throughout the training process. This procedure is time-consuming, resulting in a training duration that is roughly twice that of the other two methods. These findings indicate that the Spatial-DR method effectively enhances the generalizability of the control policy without requiring additional computational resources.
@@ III-B Performane of Control Policy
Figure 4: The curve of cumulative rewards changes with the training steps.
@@
Comparison of Regularization (Reg): To verify the effectiveness of regularization methods, we designed the following experiments. Activation Reg. This method added a regularization term to the Critic’s loss function. Discount Reg. The proposed method, it altered the discount factor $\gamma$ in the PPO algorithm to a smaller value $\gamma_{\rm reg}$. Normal. This approach does not employ regularization. Phasic Policy Gradient (PPG) [28]. A SOTA method for enhancing the generalizability of RL algorithms. It integrates auxiliary losses during the training of the PPO algorithm to improve the algorithm’s generalizability, details can be found in [28]. We plotted the cumulative reward curves for each approach during the training process, as shown in Fig. 4. It is evident that D-Reg yielded the best results, followed by A-Reg, while PPG underperformed, even falling behind Normal, which doesn’t use any regularization techniques. Taking both the effectiveness and the simplicity of algorithm implementation into account, D-Reg proves to be superior in our locomotion control tasks. It is noteworthy that PPG exhibited the poorest performance. We suspect this may be because the PPG algorithm was developed for image-based RL and might not be well suited for motion control RL with low dimensional observations.
@@
Performance of Fine-tuning: In this section, we demonstrate the effectiveness of fine-tuning, we applied two methods as follows. Standard Training. Normative training under the specific morphology without pre-training. Fine-tuning. The proposed method. The training time comparison, illustrated in Fig. 5, reveals that Standard Training begins to converge at the 2000th epoch, while Fine-tuning starts converging at just the 100th epoch. These findings validate that the proposed method can significantly accelerate the training process. For instance, when training 30 specific morphologies, Standard Training requires $30\times 6000=1.8\times 10^{5}$ epochs, whereas Fine-tuning only needs $6000+30\times 400=1.8\times 10^{4}$ epochs. The number 6000 represents the pre-training steps, while 400 indicates the fine-tuning steps tailored for specific morphologies, equivalent to 6.67% of the former. Furthermore, we conduct a performance analysis of the two methods, as illustrated in Fig. 5. For Morphology-1, both methods exhibit convergence around a value of 10, while for Morphology-2, convergence is observed around a value of 12. We calculate the mean of the last 100 reward values for both methods and assess the differences. Specifically, for Morphology-1, the difference in reward values between the two methods is 2.88%, whereas for Morphology-2, this difference is 2.08%. These findings indicate that the proposed approach can achieve performance comparable to standard training within a smaller margin of error, while requiring only one-tenth of the time.
Figure 5: Comparison of Fine-tuning and Standard Training. (a) and (b) represent the results under different morphologies.
@@ III-C Comparison with Co-design Baselines
In this section, we compare the proposed method with three co-design baselines, we first introduce these methods. Online-PCODP [6]. This method treats design parameters as privileged information and samples these parameters uniformly within their range during the training process. It assumes that near-optimal performance can be achieved using a policy conditioned on design parameters. Offline-PCODP [3]. This method collects interaction data from various morphologies using online RL and then uses offline RL to create a general policy from these data. We employed 16 distinct morphologies to gather interaction data, utilizing the TD3-BC [30] algorithm for offline training. EAT [5]. Similar to Offline-PCODP, the primary difference lies in replacing the TD3-BC method with the Decision Transformer [31]. Ours. The proposed method.
We uniformly sampled 81 different morphologies and plotted the cumulative rewards of different methods in Fig. 6. As can be seen from Fig. 6(d), our proposed method achieves the highest rewards across various morphologies. For the Online-PCODP method, as observed in Fig. 6(a), its highest reward values are concentrated under the last trained morphology. For previously encountered morphologies, it shows lower cumulative rewards. For the Offline-PCODP method, as seen in Fig. 6(b), its highest rewards are found near the 16 typical morphologies encountered during training. However, for morphologies not previously encountered, it shows lower cumulative rewards. Regarding the EAT method, as seen in Fig. 6(c), its performance appears to be poor across all morphologies. We speculate that this may be due to the original EAT method focusing on simple planar walking tasks with low-dimensional states such as body linear velocity and angular velocity. In the parkour task, terrain information is crucial but results in higher state dimensionality, potentially causing Transformer-based algorithms to fail.
Figure 6: Heatmap of co-design experiments. Each block represents the cumulative rewards achieved using the corresponding control strategy under a given morphology. For Offline-PCOD and EAT, the 16 typical morphologies are located along the diagonal of the heatmap (excluding the center point).
@@
Furthermore, we applied the aforementioned methods for co-design in both the high jump and long jump tasks. We recorded the accumulative rewards in Tab. III. It is evident that the proposed method achieved the highest values in both tasks, markedly surpassing the baseline methods. In Tab. III, we also recorded the time consumption of various methods. The introduction of the fine-tuning during the BO process has incurred some additional time compared to Online-PCODP method, however, overall, the time consumed by the proposed method is still much less than that of the Offline-PCODP and EAT methods. Taking both performance and time consumption into consideration, the advantages of the proposed method are clearly notable.
@@ III-D Parkour Performance Comparison
In this section, we compared the performance achievable by baselines with that of the proposed approach for the tasks of long jump and high jump, as illustrated in Tab. IV. The values for the baseline methods were cited from the corresponding papers. As seen in Tab. IV, the proposed method significantly enhances the performance of robots in various parkour tasks, demonstrating that, in addition to improving the motion controller’s performance, appropriate optimization of the mechanical structure can also enhance the performance of quadruped robots in extreme tasks. Furthermore, we illustrated the performance of the robots in extreme tasks before and after morphological optimization, as shown in Fig. 7. For the long jump task, it is clear that the optimized robots are equipped with longer hind legs, which can provide more robust propulsion, while shorter front legs can assist the robot in gripping the ground ahead to prevent falls. Similarly, for the high jump task, the robots evolved to have longer legs, aiding in climbing. Both experiments underscore the potential of the co-design approach in unlocking the parkour performance of robots.